(8 pts) 4) Use the graphs below to answer the questions. (4 pts each) 3 1 2 34 r -f(a) a) (g o f)(1) c) fo g)(4)
Answer #1

(a) We have f(1) = 2 and (2) = 3.

Therefore we have

(gof ) ( 1 )=g(/1))=g(2)-3.

(b) We have g(1)=3 and f (3) 3

Therefore

fog) (1g(1)- f(3)-3.

(c) We have (4) = 4 and f(4)=0.

Therefore

(fog)(4)=f(g(4))=f(4) =0